# Multi-Scale Spatial-Temporal Hypergraph Network with Lead-Lag Structures for Stock Time Series Forecasting

### v1

@@ 3.4 Hyperedge-based Moving Aggregation
After delineating different time series scales with Spatial-Temporal hypergraphs $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots\mathcal{G}^{S}\}$, we further consider the lead-lag relationships between industries under each scale, to analyze the complex correlations in financial markets—see Figure 4 in Appendix B.4.
@@
where $\mathcal{P}^{i}\in\mathbb{R}^{K\times(T^{i}-k^{i}+1)\times k^{i}\times d}$ denotes the patches of $E^{i}$, containing local information with the size of $k^{i}$. From the perspective of causality, considering the relationship between the first few points and the last point within the window, we can construct the relationship matrix of the leading part and the lagging part. We then mine the lead-lag relationships inside each patch by considering the correlations between $\mathcal{P}^{i}_{before}$ and $\mathcal{P}^{i}_{last}$:
$\mathcal{P}^{i}_{before}=\text{Reshape}(\mathcal{P})^{i}\in\mathbb{R}^{[K\times(T^{i}-k^{i}+1)]\times k^{i}\times d},$ (10)
$\mathcal{P}^{i}_{last}=\text{Reshape}(\mathcal{P})^{i}_{-1}\in\mathbb{R}^{(K\times 1)\times k^{i}\times d},$ (11)
@@
$\tilde{A}^{i}=\mathcal{P}^{i}_{before}\otimes\mathcal{P}^{i}_{last}\in\mathbb{R}^{[K\times(T^{i}-k^{i}+1)]\times K\times k^{i}},$ (12)
$A^{i}=SoftMax(\tilde{A}^{i},\text{dim}=2),~~\tilde{E}^{i}=A^{i}\otimes\mathcal{P}^{i}_{before}\in\mathbb{R}^{K\times k^{i}\times d},$ (13)
$\hat{E}^{i}=\text{MLP}^{i}(\tilde{E}^{i})+E^{i},~~\text{MLP}^{i}:=\mathbb{R}^{k^{i}}\to\mathbb{R}^{T^{i}},~~\mathcal{V}^{e^{i}}=\hat{E}^{i}\otimes\mathcal{H}\in\mathbb{R}^{N\times T^{i}\times d},$ (14)
where $\tilde{A}^{i}$ considers the correlations among the leading parts and lagging parts, $A^{i}$ probabilityizes $\tilde{A}^{i}$ to control weightsuming, $\hat{E}^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ denotes the embedding of hyperedges integrated with lead-lag relationships among different industries. $\mathcal{V}^{e^{i}}$ is the updated nodes of $i$-th scale integrated with the lead-lag information extracted on hyperedges—see Figure 4 in Appendix B.4.
@@ 3.5 Hyperedge-based Multi-scale Fusion
$\hat{\mathcal{E}}^{i}=\text{Linear}^{i}(\hat{E}^{i})\in\mathbb{R}^{K\times T\times d},~~\text{Linear}^{i}=\mathbb{R}^{T^{i}}\to\mathbb{R}^{T},$ (15)
$\hat{\mathcal{E}}=\{\hat{\mathcal{E}}^{1},\hat{\mathcal{E}}^{2},\cdots,\hat{\mathcal{E}}^{S}\}\in\mathbb{R}^{S\times K\times T\times d}$ (16)
Then, we construct adaptive adjacency matrices for these hyperedges of different scales to measure the correlations between different industries at various information granularities. We employ metric learning for this task and use a learnable Mahalanobis distance to automatically build relationships between different industries in the hidden space:
$\tilde{\mathcal{E}}=\text{Flatten}(\hat{\mathcal{E}})\in\mathbb{R}^{(S\times K)\times T\times d},~~D_{ij}=(\tilde{\mathcal{E}}_{i}-\tilde{\mathcal{E}}_{j})^{T}\otimes Q\otimes(\tilde{\mathcal{E}}_{i}-\tilde{\mathcal{E}}_{j}),$ (17)
$C_{ij}=\begin{cases}\frac{1}{D_{ij}}&\text{if }i\neq j\\ 0&\text{if }i=j\end{cases},~~B=\text{SoftMax}(C,\text{dim}=0)$ (18)
where $D,C,B\in\mathbb{R}^{(S\times K)\times(S\times K)\times T\times d}$. $Q\in\mathbb{R}^{d\times d}$ is a learnable semi-positive definite matrix, and can be practically constructed by $Q=L^{T}\cdot L$. Subsequently, we construct interactions between hyperedges at different scales using message passing and further update the representation of hyperedges.
$\tilde{B}=\mathcal{D}^{-1}\otimes B\otimes\mathcal{D},~~\mathcal{E}^{\prime}=\tilde{B}\otimes\tilde{\mathcal{E}}\otimes W^{e},$ (19)
where $\mathcal{D}\in\mathbb{R}^{(S\times K)\times T\times d}$ is the degree matrix, $W^{e}\in\mathbb{R}^{d\times d}$ is the learnable parameters in the graph message passing mechanism. After integrating the lead-lag relationships between industries on the hyperedges and fusing the hidden relationships between different scales, the representation of the hyperedges $\mathcal{E}^{\prime}$ is adaptively updated. We then continue to complete the message passing from hyperedges to nodes—see Figure 4 in Appendix B.4, thereby feeding back the complex correlations within the industry to specific agents and completing the update of the representation of agent nodes:
$\mathcal{V}$ $=\mathcal{E}^{\prime}\otimes\mathcal{H}\in\mathbb{R}^{(S\times N)\times T\times d}$ (20)
@@ 3.6 Predictor
After completing the complex representation extraction using a hypergraph-based structure, we employ an MLP-based predictor and incorporate residual connection modules to forecast financial indicators across each scale $s$.
$\hat{\mathcal{X}}^{s}=\text{MLP}_{s,1}(\mathcal{X}^{s})\in\mathbb{R}^{N\times 1},~~\tilde{\mathcal{V}}^{s}=\text{MLP}_{s,2}(\mathcal{V}^{s})\in\mathbb{R}^{N\times 1},$ (21)
$\hat{Y}^{s}=\hat{\mathcal{X}}^{s}+\tilde{\mathcal{V}}^{s},~~\hat{Y}=\sum_{s=1}^{S}\hat{Y}^{s},$ (22)
where $X^{s}\in\mathbb{R}^{(N\times T^{s}\times d)}$ represents the initial node embedding in Equation 5 that does not consider the lead-lag relationship and multi-scale interaction effects. However, $\mathcal{V}^{s}\in\mathbb{R}^{(N\times T\times d)}$ is the spatio-temporal hypernode embedding in Equation 20 that incorporates lead-lag correlation and multi-scale information. Finally, we combine the prediction results at different scales to obtain the final prediction result.
@@ 3.7 Optimization Objectives
Following StockMixer (Fan & Shen, 2024), we use the return ratio of a stock as the groundtruth. We use a combination of a pointwise regression and pairwise ranking-aware loss to minimize the MSE between the predicted and actual 1-day return ratios, while maintaining the relative order of top-ranked stocks with higher expected return for investment as follows.
[…]
$\mathcal{L}=\mathcal{L}_{\mathit{MSE}}+\alpha\mathcal{L}_{Rank},$ (24)
where $\hat{r}^{t}$ and ${r}^{t}$ are the predicted and actual ranking scores, respectively, and $\alpha$ is a weight parameter.
@@ 4.1 Experimental Setup
Datasets. We evaluate our approach using three real-world datasets from the US stock market. These datasets contain relatively complete stock-industry relationships or Wiki-based company relations. The NASDAQ and NYSE datasets (Feng et al., 2019b) include EOD data from 01/02/2013 to 12/08/2017, filtered from the respective markets. Abnormal patterns and penny stocks were removed, while preserving their representative properties, with NASDAQ being more volatile and NYSE more stable. The S&P500 dataset (Huynh et al., 2023) gathers historical price data and industry information for the companies listed in the S&P500 index from the Yahoo Finance database, covering the period from 01/04/2016 to 05/25/2022. The datasets statistics are provided in Table 4.
Baselines. We conduct a comparative analysis between our proposed Hermes and a range of baseline methods, which include prominent approaches in both machine learning and deep learning methods for STSF. Given that existing research has demonstrated the limited effectiveness of traditional statistical methods (Fan & Shen, 2024; Li et al., 2024), we exclude them from our comparative evaluation to focus on more competitive approaches.
Table 1: Comparison results on public stock market datasets (measured by t-test with p-value < 0.01). The methods for comparison are mainly divided into four types: RNN (Recurrent Neural Network), GNN (Graph Neural Network), HGNN (HyperGraph Neural Network), and MLP (Multi-Layer Perceptron). Bold indicates the best (SOTA) results.
Model NASDAQ NYSE S&P500
IC RIC prec@N SR IC RIC prec@N SR IC RIC prec@N SR
RNN LSTM 0.032 0.354 0.514 0.892 0.024 0.256 0.512 0.857 0.031 0.186 0.531 1.332
ALSTM 0.035 0.371 0.522 0.941 0.023 0.276 0.519 0.764 0.029 0.181 0.532 1.298
GNN RGCN 0.034 0.382 0.516 1.054 0.025 0.275 0.517 0.932 0.028 0.175 0.528 1.359
GAT 0.035 0.377 0.530 1.233 0.025 0.297 0.521 1.070 0.034 0.191 0.541 1.484
RSR-I 0.038 0.398 0.531 1.238 0.026 0.284 0.519 0.098 0.033 0.200 0.542 1.437
MLP Linear 0.019 0.188 0.505 0.517 0.015 0.163 0.497 0.625 0.016 0.156 0.520 0.674
StockMixer 0.043 0.501 0.545 1.465 0.029 0.351 0.539 1.454 0.041 0.262 0.551 1.586
HGNN STHAN-SR 0.039 0.451 0.543 1.416 0.029 0.344 0.542 1.228 0.037 0.227 0.549 1.533
ESTIMATE 0.037 0.444 0.539 1.307 0.030 0.327 0.536 1.115 0.035 0.241 0.553 1.547
Hermes (Ours) 0.044 0.538 0.535 2.161 0.032 0.466 0.549 1.655 0.050 0.334 0.539 2.247
@@ 4.2 Overall Performance
We present a detailed comparison of our approach (Hermes) with four types of baseline methods in Table 1. We have the following observations: 1) For univariate methods, whether it is LSTM or ALSTM, their performance is inferior to that of other hybrid architectures. This strongly highlights the necessity and importance of capturing the complex industry and stock relationships in financial markets. 2) For GNNs algorithms such as RGCN, GAT, and RSR-I, their performance shows a significant improvement over RNNs, indicating that GNNs have strong stock relationship modeling capabilities. However, compared to HGNNs algorithms like STHAN-SR and ESTIMATE, GNNs perform slightly worse. This suggests that in the context of the financial market industry, using hypergraph models to unify the relationships of stocks within the industry plays a crucial role in enhancing forecasting performance. 3) The simple linear model fails due to insufficient inductive bias. While StockMixer strikes a balance between the simplicity of MLP models and the excellent performance of hybrid networks, it achieves the second-best results across most metrics. However, because MLP-based models inherently have smaller model capacities, the issue of insufficient inductive bias in StockMixer becomes more apparent when dealing with larger datasets. 4) Compared to other hypergraph algorithms that consider stock relationships within industries (such as STHAN-SR and ESTIMATE), the Hermes algorithm we propose shows significant performance improvement. This clearly demonstrates that, based on the consideration of stock relationships within industries, incorporating inter-industry lead-lag relationships and multi-scale relationships plays a crucial role in enhancing forecasting capability.

### v2

@@ 3.4 Hyperedge-based Moving Aggregation
After delineating different time series scales with Spatial-Temporal hypergraphs $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots\mathcal{G}^{S}\}$, we further consider the lead-lag relationships between industries under each scale to analyze the complex correlations in financial markets—see Figure 5 in Appendix B.4.
@@
where $\mathcal{P}^{i}\in\mathbb{R}^{K\times(T^{i}-k^{i}+1)\times k^{i}\times d}$ denotes the patches of $E^{i}$, containing local information with the size of $k^{i}$. From the perspective of causality, considering the relationship between the entire window and its last point, we can construct the relationship matrix of the leading part and the lagging part. We then mine the lead-lag relationships inside each patch by considering the correlations between $\mathcal{P}^{i}_{before}$ and $\mathcal{P}^{i}_{last}$:
$\mathcal{P}^{i}_{before}=\text{Reshape}(\mathcal{P})^{i}\in\mathbb{R}^{(T^{i}-k^{i}+1)\times[K\times k^{i}]\times d},$ (14)
$\mathcal{P}^{i}_{last}=\text{Reshape}(\mathcal{P})^{i}_{-1}\in\mathbb{R}^{(T^{i}-k^{i}+1)\times[K\times 1]\times d}.$ (15)
@@
$\tilde{A}^{i}=\mathcal{P}^{i}_{before}\otimes\mathcal{P}^{i}_{last}\in\mathbb{R}^{(T^{i}-k^{i}+1)\times[K\times k^{i}]\times K},$ (16)
$A^{i}=SoftMax(\tilde{A}^{i},\text{dim}=1),$ (17)
$\tilde{E}^{i}=A^{i}\otimes\mathcal{P}^{i}_{before}\in\mathbb{R}^{K\times(T^{i}-k^{i}+1)\times d},$ (18)
$\hat{E}^{i}\in\mathbb{R}^{K\times T^{i}\times d}=\begin{cases}E^{i},&t<k^{i}\\ {\tilde{E}^{i}},&k^{i}\leq t\leq T^{i}\end{cases},$ (19)
where $\tilde{A}^{i}$ considers the correlations among the leading parts and lagging parts, $A^{i}$ probabilityizes $\tilde{A}^{i}$ to control weightsuming, $\hat{E}^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ denotes the embedding of hyperedges integrated with lead-lag relationships among different industries. It is particularly important to note that when $t<k^{i}$, no transformation is performed; when $k^{i}\leq t\leq T^{i}$, the result incorporating lead-lag interactions is used as the new hyperedges.
@@ 3.5 Hyperedge-based Multi-scale Fusion
$\hat{\mathcal{E}}^{i}=\text{Linear}^{i}(\hat{E}^{i})\in\mathbb{R}^{K\times T\times d},\text{Linear}^{i}=\mathbb{R}^{T^{i}}\to\mathbb{R}^{T},$ (20)
$\hat{\mathcal{E}}=\{\hat{\mathcal{E}}^{1},\hat{\mathcal{E}}^{2},\cdots,\hat{\mathcal{E}}^{S}\}\in\mathbb{R}^{S\times K\times T\times d}.$ (21)
Then, we construct adaptive adjacency matrices for these hyperedges of different scales to measure the correlations between different industries at various information granularities. We employ metric learning for this task and use a learnable Mahalanobis distance Goldberger et al. (2004); Qiu et al. (2025b) to automatically build relationships between different industries in the hidden space. Since the operations across the time dimension $T$ are independent and can be executed in parallel, we omit this dimension in the following formulations for notational simplicity.
$\tilde{\mathcal{E}}=\text{Flatten}(\hat{\mathcal{E}})\in\mathbb{R}^{(S\times K)\times d},$ (22)
$D_{ij}=(\tilde{\mathcal{E}}_{i}-\tilde{\mathcal{E}}_{j})\cdot Q\cdot(\tilde{\mathcal{E}}_{i}-\tilde{\mathcal{E}}_{j})^{T},$ (23)
$C_{ij}=\begin{cases}\frac{1}{D_{ij}}&\text{if }i\neq j\\ 0&\text{if }i=j\end{cases},~~B=\text{SoftMax}(C,\text{dim}=0),$ (24)
where $D,C,B\in\mathbb{R}^{(S\times K)\times(S\times K)}$. $Q\in\mathbb{R}^{d\times d}$ is a learnable semi-positive definite matrix, and can be practically constructed by $Q=A^{T}\cdot A$, where A is also a learnable matrix. Subsequently, we construct interactions between hyperedges at different scales using message passing and further update the representation of hyperedges.
$\mathcal{E}^{\prime}=B\cdot\tilde{\mathcal{E}}\cdot W^{e}\in\mathbb{R}^{(S\times K)\times d},$ (25)
where $W^{e}\in\mathbb{R}^{d\times d}$ is the learnable parameters in the graph message passing mechanism. After integrating the lead-lag relationships between industries on the hyperedges and fusing the hidden relationships between different scales, the representation of the hyperedges $\mathcal{E}^{\prime}$ is adaptively updated. We then continue to complete the message passing from hyperedges to nodes—see Figure 5 in Appendix B.4, thereby feeding back the complex correlations within the industry to specific stocks and completing the update of the representation of stock nodes. Finally, we restore the omitted $T$ dimension:
$\mathcal{V}$ $=\mathcal{E}^{\prime}\otimes\mathcal{H}\in\mathbb{R}^{(S\times N)\times T\times d}.$ (26)
@@ 3.6 Predictor
After completing the complex representation extraction using a hypergraph-based structure, we employ an MLP-based predictor and incorporate residual connection modules to forecast financial indicators across each scale $i$.
$\hat{\mathcal{X}}^{i}=\text{MLP}^{i}_{1}(\mathcal{X}^{i})\in\mathbb{R}^{N\times 1},$ (27)
$\tilde{\mathcal{V}}^{i}=\text{MLP}^{i}_{2}(\mathcal{V}^{i})\in\mathbb{R}^{N\times 1},$ (28)
$\hat{Y}^{i}=\hat{\mathcal{X}}^{i}+\tilde{\mathcal{V}}^{i},~~\hat{Y}=\sum_{i=1}^{S}\hat{Y}^{i},$ (29)
where $\mathcal{X}^{i}\in\mathbb{R}^{(N\times T^{i}\times d)}$ represents the initial node embedding in Equation 6 that does not consider the lead-lag relationship and multi-scale interaction effects. However, $\mathcal{V}^{i}\in\mathbb{R}^{(N\times T\times d)}$ is the spatio-temporal hypernode embedding in Equation 26 that incorporates lead-lag correlation and multi-scale information. Finally, we combine the prediction results at different scales to obtain the final prediction result.
Models HGNN Transformer MLP GNN RNN
Hermes EST. STHAN. ASHyper iTrans. PatchTST DUET SMixer. DLin. TMixer. RSR-I GAT RGCN ALSTM LSTM
NASDAQ IC 0.044 0.037 0.039 0.041 0.038 0.039 0.039 0.043 0.031 0.041 0.038 0.035 0.034 0.035 0.032
RIC 0.538 0.444 0.451 0.508 0.420 0.496 0.494 0.501 0.453 0.532 0.398 0.377 0.382 0.371 0.354
SR 2.161 1.307 1.416 1.157 1.877 1.261 1.751 1.465 1.778 1.177 1.238 1.233 1.054 0.941 0.892
NYSE IC 0.032 0.030 0.029 0.025 0.023 0.029 0.024 0.029 0.026 0.031 0.026 0.025 0.025 0.023 0.024
RIC 0.466 0.327 0.344 0.250 0.303 0.284 0.327 0.351 0.277 0.314 0.284 0.297 0.275 0.276 0.256
SR 1.655 1.115 1.228 1.208 1.520 0.898 1.361 1.454 1.526 1.556 0.098 1.070 0.932 0.764 0.857
S&P500 IC 0.050 0.035 0.037 0.046 0.048 0.046 0.049 0.041 0.043 0.040 0.033 0.034 0.028 0.029 0.031
RIC 0.334 0.241 0.227 0.268 0.242 0.249 0.275 0.262 0.214 0.253 0.200 0.191 0.175 0.181 0.186
SR 2.247 1.547 1.533 2.148 1.326 1.402 1.982 1.586 1.557 2.170 1.437 1.484 1.359 1.298 1.332
Model abbreviations: EST. (ESTIMATE), STHAN. (STHAN-SR), iTrans. (iTransformer), SMX (StockMixer), DLin. (DLinear), and TMX (TimeMixer).
Table 1: Comparison results on public stock market datasets (measured by t-test with p-value < 0.01). The methods for comparison are mainly divided into five types: HGNN (HyperGraph Neural Network), Transformer, MLP (Multi-Layer Perceptron), GNN (Graph Neural Network), and RNN (Recurrent Neural Network). Bold indicates the best (SOTA) results.
@@ 3.7 Optimization Objectives
Following StockMixer Fan and Shen (2024), We use the 1-day return ratio of a stock as the ground-truth rather than the normalized price used in previous work using a combination of a pointwise regression and pairwise ranking-aware loss to minimize the MSE between the predicted and actual return ratios while maintaining the relative order of top ranked stocks with higher expected return for investment as:
$L=L_{\text{MSE}}+\alpha\displaystyle\sum_{i=1}^{N}\sum_{j=1}^{N}\max(0,-(\hat{r}_{i}^{t}-\hat{r}_{j}^{t})(r_{i}^{t}-r_{j}^{t})),$ (30)
where $\hat{r}^{t}$ and ${r}^{t}$ are the predicted and actual ranking scores, respectively, and $\alpha$ is a weighting parameter.
@@ 4.1 Experimental Setup
Datasets. We evaluate our approach using three real-world datasets from the US stock market. These datasets contain relatively complete stock-industry relationships or Wiki-based company relations. The NASDAQ and NYSE datasets Feng et al. (2019b) include EOD data from 01/02/2013 to 12/08/2017, filtered from the respective markets. Abnormal patterns and penny stocks were removed, while preserving their representative properties, with NASDAQ being more volatile and NYSE more stable. The S&P500 dataset Huynh et al. (2023) gathers historical price data and industry information for the companies listed in the S&P500 index from the Yahoo Finance database, covering the period from 01/04/2016 to 05/25/2022. The datasets statistics are provided in Table 4.
Baselines. We conduct a comparative analysis between our proposed Hermes and a range of baseline methods, which include prominent approaches in both machine learning and deep learning methods for STSF. Given that existing research has demonstrated the limited effectiveness of traditional statistical methods Fan and Shen (2024); Li et al. (2024), we exclude them from our comparative evaluation to focus on more competitive approaches.
Implementation Details. For fair comparison, all samples are generated by moving a 16 time steps lookback window. All experiments with Hermes are conducted using PyTorch Paszke et al. (2019) in Python 3.9, and are executed on a server featuring an Intel(R) Xeon(R) Platinum 8358 CPU and an NVIDIA Tesla-A800 GPU. We use ADAM Kingma and Ba (2015) with an initial learning rate of 5e-3. The “Drop Last” issue is reported by several researchers Qiu et al. (2024, 2025a). That is, in some previous works evaluating the model on test set with drop-last=True setting may cause additional errors related to test batch size. In our experiment, to ensure fair comparison in the future, we set the drop last to False for all baselines to avoid this issue.
Detailed descriptions of the experimental datasets, baselines, and evaluation metrics are provided in the Appendix A.1.
Ablation NASDAQ NYSE S&P500
Model Component IC ICIR SR IC ICIR SR IC ICIR SR
w/o Multi-Scale Fusion 0.038 0.464 1.374 0.028 0.421 1.354 0.031 0.248 1.517
w/o Total Multi-Scale 0.030 0.422 0.975 0.023 0.347 1.199 0.032 0.213 1.075
w/o Lead-Lag 0.039 0.472 1.451 0.028 0.420 1.225 0.037 0.201 1.628
w/o Skip-Connection 0.041 0.441 1.655 0.029 0.345 1.449 0.043 0.252 1.653
Only Initial Node Embedding 0.019 0.174 0.695 0.016 0.113 0.449 0.018 0.122 0.659
Hermes (Ours) 0.044 0.538 2.161 0.032 0.466 1.655 0.050 0.334 2.247
Dataset NASDAQ NYSE S&P500
Property STHAN. EST. ASHyper Hermes STHAN. EST. ASHyper Hermes STHAN. EST. ASHyper Hermes
Training Time 257 ms 1,267 ms 288 ms 215 ms 384 ms 1,462 ms 378 ms 282 ms 297 ms 1,317 ms 314 ms 211 ms
Inference Time 52 ms 163 ms 58 ms 42 ms 73 ms 241 ms 71 ms 61 ms 32 ms 109 ms 36 ms 25 ms
Memory Usage 2.15 GB 9.62 GB 1.99 GB 0.98 GB 2.82 GB 10.25 GB 2.17 GB 1.14 GB 1.23 GB 4.52 GB 1.51 GB 0.54 GB
Table 3: Efficiency analysis on public stock datasets, concerning the training time, inference time (measured in milliseconds per batch under the condition of a batch size being 1), and memory usage (quantified in GB) of different HGNN-based (HyperGraph Neural Network) models. Model abbreviations: EST. (ESTIMATE) and STHAN. (STHAN-SR).
@@ 4.2 Overall Performance
We conduct a comprehensive comparison of Hermes against five categories of baseline methods (HGNN, Transformer, MLP, GNN, and RNN) across three representative stock market datasets: NASDAQ, NYSE, and S&P 500. As demonstrated in Table 1. We have the following observations: 1) Hermes consistently achieves State-of-the-Art (SOTA) performance across all datasets and evaluation metrics. 2) For univariate methods, whether it is LSTM or ALSTM, their performance is inferior to that of other hybrid architectures. This strongly highlights the necessity and importance of capturing the complex industry and stock relationships in financial markets. 3) For GNNs algorithms such as RGCN, GAT, and RSR-I, their performance shows a significant improvement over RNNs, indicating that GNNs have strong stock relationship modeling capabilities. However, compared to HGNNs algorithms like STHAN-SR, ESTIMATE, and ASHyper, GNNs perform slightly worse. This suggests that in the context of the financial market industry, using hypergraph models to unify the relationships of stocks within the industry plays a crucial role in enhancing forecasting performance. 4) Compared to other hypergraph algorithms that consider stock relationships within industries (such as STHAN-SR, ESTIMATE, and ASHyper), the Hermes algorithm we propose shows significant performance improvement. This clearly demonstrates that, based on the consideration of stock relationships within industries, incorporating inter-industry lead-lag relationships and multi-scale relationships plays a crucial role in enhancing forecasting performance.